# Superconductivity in three-dimensional interacting doped topological insulators

### v1

@@ Abstract
To gain insights into the nature of the superconducting ground states in the $(U_{1},U_{2})$ plane, it is worth focusing on the $U_{1}$ and $U_{2}$ axes separately. As the strength of repulsive $U_{1}$ is increased beyond its critical value, $g^{s}_{0,1,3}\to+\infty$ and $g^{s}_{2}\to-\infty$, becoming effectively repulsive and attractive, respectively. From the unbiased RG procedure it can be shown that bare repulsive $g^{s}_{0,1,3}$ and bare attractive $g^{s}_{2}$ favor $s$-wave pairing, as also found by tuning solely $U_{1}$. On the other hand, when the strength of $U_{2}$ is tuned beyond the critical one, $g^{s}_{0,1,2}\to+\infty$ while $g^{s}_{3}\to-\infty$. RG flow with bare repulsive $g^{s}_{0,1}$ favors $s$-wave pairing, whereas bare repulsive (attractive) $g^{s}_{2}$ ($g^{s}_{3}$) favors pseudoscalar $p$-wave pairing. When the strength of only $U_{2}$ is tuned, the degree of divergence for $g^{s}_{2,3}$ is larger than that for the $g^{s}_{0,1}$, causing condensation of odd-parity pseudoscalar pairing. These observation can be further substantiated from recently proposed “selection rules” (SRs) between the dominant interaction and the resulting ordered state(s) [44, 45, 46], which we discuss next.
@@
Selection rule. We focus on a single quartic term of the form $g_{Q}\sum_{i}(\Psi^{\dagger}_{\rm Nam}Q_{i}\Psi_{\rm Nam})^{2}$ and a fermion bilinear of the form $\sum_{i}\Psi^{\dagger}_{\rm Nam}R_{i}\Psi_{\rm Nam}$, chosen from the sets of all the symmetry-allowed four-fermion interactions [Eq. (2)] and order parameters [Eqs. (7)-(9)], respectively. Here $Q_{i}$ and $R_{i}$ are Hermitian matrices, expressed in the announced Nambu basis. Let $A_{M}$ be the number of anticommuting matrix pairs between the four-fermion interaction and order parameter terms. For concreteness, we now consider all the interactions to be repulsive ($g_{Q}>0$). Then among all the available ordered phases, the interaction channel $g_{Q}$ maximally boosts the nucleation of the ones for which (SR1) $Q_{i}\equiv R_{i}$ or (SR2) $A_{M}$ is maximal. It is also conceivable for a given repulsive interaction to support excitonic (for zero and low $\mu$) and superconducting (for moderate $\mu$) states, for example, described by the fermion bilinears $\sum^{K}_{i=1}\Psi^{\dagger}_{\rm Nam}R^{\rm exc}_{i}\Psi_{\rm Nam}$ and $\sum^{L}_{i=1}\Psi^{\dagger}_{\rm Nam}R^{\rm pair}_{i}\Psi_{\rm Nam}$, respectively, transforming as vectors under O($K$) and O($L$) rotations, constituted by $K$ and $L$ number of mutually anticommuting matrices. Then these two ordered phases can appear next to each other when, for example, $\mu$ is tuned in the system, if (SR3) the corresponding vector order parameters form composite vectors that transform under the O($N$) rotations, where $K,L<N\leq K+L$. These SRs have been successfully employed for twisted bilayer graphene [44], 3D (un)doped Luttinger materials [45], and (un)doped monolayer and bilayer graphene [46].
@@
We exemplify such a seemingly mathematical formulation by considering a specific (repulsive) interaction $g^{s}_{2}$ in doped Dirac insulator. For low chemical doping, it supports nucleation of the excitonic pseudoscalar mass, transforming as an O(1) or a $Z_{2}$ vector, following the SR1. In the moderate doping regime, the same interaction favors condensation of the pseudoscalar pairing, transforming as an O(2) vector, following the SR2. In the $(\mu,t)$ plane of the phase diagram with a fixed $g^{s}_{2}>0$, shown in the SM [42], these ordered phases live next to each other, as they form a composite O(3) vector [47], following the SR3. These outcomes qualitatively justify the appearance of the pseudoscalar pairing along the $U_{2}$ axis [Fig. 1(b)-(d)]. Similar conclusions hold when we tune any one of the eight local four-fermion interactions [Eq. (2)]. The results are summarized in Table 2, and the corresponding cuts of the phase diagrams in the $(\mu,t)$ plane are displayed in the SM [42].
@@
Summary & discussion. Here we formulate an unbiased and controlled RG framework to capture dominant superconducting instabilities of 3D doped interacting Dirac insulators (topological and trivial), featuring an intriguing confluence of topological $p$-wave and trivial $s$-wave pairings, triggered by on-site Hubbard repulsions. Our approach is sufficiently general to address similar phenomenon in doped TIs belonging to any Altland-Zirnbauer symmetry class in any dimension $d>1$, as well as in doped crystalline TIs, harboring FSs near multiple high symmetry points in the Brillouin zone, connected by discrete crystal symmetries, and in doped higher-order TIs and topological Dirac and Weyl semimetals. The present analysis can be generalized to encompass longer, but finite range Coulomb repulsions as well as layered topological materials possessing tetragonal symmetry, for example, for which $w_{x}=w_{y}\neq w_{z}$, where $w=v,b,a$. These generalizations may allow appearance of other paired states (such as the nematic one), tabulated in Table 1, in their global phase diagram. The current work lays the theoretical foundation for such fascinating future research directions.
@@
Acknowledgments. A.S. is grateful for financial support from the Swiss National Science Foundation (SNSF) through Division II (No. 184739). B.R. was supported by NSF CAREER Grant No. DMR-2238679, and thankful to Vladimir Juričić and Sanjib Kumar Das for critical reading of the manuscript.

### v2

@@ Abstract
To gain insights into the nature of the superconducting ground states in the $(U_{1},U_{2})$ plane, it is worth focusing on the $U_{1}$ and $U_{2}$ axes separately. As the strength of repulsive $U_{1}$ is increased beyond its critical value, $g^{\rm s}_{0,1,3}\to+\infty$ ($g^{\rm s}_{2}\to-\infty$), becoming effectively repulsive (attractive). From the unbiased RG procedure it can be shown that bare repulsive $g^{\rm s}_{0,1,3}$ and bare attractive $g^{\rm s}_{2}$ favor $s$-wave pairing, as also found by tuning solely $U_{1}$. On the other hand, when the strength of $U_{2}$ is tuned beyond the critical one, $g^{\rm s}_{0,1,2}\to+\infty$ while $g^{\rm s}_{3}\to-\infty$. RG flow with bare repulsive $g^{\rm s}_{0,1}$ favors $s$-wave pairing, whereas bare repulsive (attractive) $g^{\rm s}_{2}$ ($g^{\rm s}_{3}$) favors pseudoscalar $p$-wave pairing. When the strength of only $U_{2}$ is tuned, the degree of divergence for $g^{\rm s}_{2,3}$ is larger than that for the $g^{\rm s}_{0,1}$, causing condensation of odd-parity pseudoscalar pairing. These observation can be further substantiated from recently proposed “selection rules” (SRs) between the dominant interaction and the resulting ordered state(s) [44, 45, 46], which we discuss next.
@@
Selection rules. We consider a single quartic term and a fermion bilinear, $g_{Q}\sum_{i}(\Psi^{\dagger}_{\rm Nam}Q_{i}\Psi_{\rm Nam})^{2}$ and $\sum_{i}\Psi^{\dagger}_{\rm Nam}R_{i}\Psi_{\rm Nam}$, respectively, chosen from the sets of all the symmetry-allowed four-fermion interactions [Eq. (2)] and order parameters [Eqs. (7)-(9)]. Here $Q_{i}$ and $R_{i}$ are Hermitian matrices, expressed in the announced Nambu basis. Let $A_{M}$ be the number of anticommuting matrix pairs between the four-fermion interaction and order parameter terms. For concreteness, we consider all the interactions to be repulsive ($g_{Q}>0$). Then among all the available ordered phases, the interaction channel $g_{Q}$ maximally boosts the nucleation of the ones for which $Q_{i}\equiv R_{i}$ (SR1) or $A_{M}$ is maximal (SR2). It is also conceivable for a given repulsive interaction to support excitonic (for zero and low $\mu$) and superconducting (for moderate $\mu$) states, described by the fermion bilinears $\sum^{K}_{i=1}\Psi^{\dagger}_{\rm Nam}R^{\rm exc}_{i}\Psi_{\rm Nam}$ and $\sum^{L}_{i=1}\Psi^{\dagger}_{\rm Nam}R^{\rm pair}_{i}\Psi_{\rm Nam}$, respectively, transforming as vectors under $O(K)$ and $O(L)$ rotations, constituted by $K$ and $L$ number of mutually anticommuting matrices, respectively. Then these two ordered phases can appear next to each other when, for example, $\mu$ is tuned in the system, if the corresponding vector order parameters form composite vectors that transform under the $O(N)$ rotations, where $K,L<N\leq K+L$ (SR3). These SRs have been successfully employed for twisted bilayer graphene [44], 3D (un)doped Luttinger materials [45], and (un)doped monolayer and bilayer graphene [46].
@@
We exemplify such a seemingly mathematical formulation by considering a specific (repulsive) interaction $g^{\rm s}_{2}$ in doped Dirac insulator. For low chemical doping, it supports nucleation of the excitonic pseudoscalar mass, transforming as an O(1) or a $Z_{2}$ vector, following the SR1. In the moderate doping regime, the same interaction favors condensation of the pseudoscalar pairing, transforming as an O(2) vector, following the SR2. In the $(\mu,t)$ plane of the phase diagram with a fixed $g^{\rm s}_{2}>0$, shown in the SM [42], these ordered phases live next to each other, as they form a composite O(3) vector [47], following the SR3. These outcomes qualitatively justify the appearance of the pseudoscalar pairing along the $U_{2}$ axis [Figs. 1(b)-(d)]. Similar conclusions hold when we tune any one of the eight local four-fermion interactions [Eq. (2)]. The results are summarized in Table 2, and the corresponding cuts of the phase diagrams in the $(\mu,t)$ plane for fixed values of $g^{{\rm s}/{\rm t}}_{\beta}>0$ are shown in the SM [42].
Notice that emergence of the superconductivity from purely repulsive electron-electron interactions in doped TIs, featuring a FS, follows the spirit of the Kohn-Luttinger mechanism [48, 49, 50]. The presence of a FS suppresses the interband scattering, responsible for particle-hole or excitonic orderings, while the intraband scattering remains unaffected yielding superconducting ground states, shown in the phase diagram of the Hubbard model in Fig. 1(b), for example. Nonetheless, as summarized in Table 2, repulsive electronic interactions also accommodate various excitonic orders, mentioned in Table 1, especially in the absence of an underlying FS [42]. Here we support these generic outcome from a controlled and unbiased RG analysis. As there is no Fermi-surface nesting in doped TIs, no charge- or spin-density-wave ordering can develop via nesting, and all the particle-hole or excitonic orderings result from inter-band scattering, which gets suppressed by chemical doping.
@@
Summary and discussion. Here, we formulate an unbiased and controlled RG framework to capture dominant superconducting instabilities of 3D doped interacting Dirac insulators (topological and trivial), featuring an intriguing confluence of topological $p$-wave and trivial $s$-wave pairings, triggered by on-site Hubbard repulsions. Our approach is sufficiently general to address similar phenomenon in doped TIs belonging to any Altland-Zirnbauer symmetry class in any dimension $d>1$, with special emphasis on two-dimensional doped TIs [51, 52, 53], as well as in doped crystalline TIs, harboring FSs near multiple high symmetry points in the Brillouin zone, connected by discrete crystal symmetries, and in doped higher-order TIs and topological Dirac and Weyl semimetals. The present analysis can be generalized to encompass longer, but finite-range Coulomb repulsions as well as layered topological materials possessing tetragonal symmetry, for which $w_{x}=w_{y}\neq w_{z}$, where $w=v,b,a$. These generalizations may allow the appearance of other competing paired states (such as the nematic one), tabulated in Table 1, in their global phase diagram. The current work lays the theoretical foundation for such fascinating future research directions.
@@
Acknowledgments. A.S. is grateful for financial support from the Swiss National Science Foundation (SNSF) through Division II (No. 184739). B.R. was supported by NSF CAREER Grant No. DMR-2238679, and is thankful to Vladimir Juričić and Sanjib Kumar Das for a critical reading of the manuscript.